George is 10 years older than his brother Sam find expression for
a) The sum for their ages
b) Sum for their ages in 5 years time
c) The product for their ages

Answer by checkley77(12844)

LET GEORGE=X
LET SAM=X-10
X+X-10= SUM.
(X+5)+(X-5)= SUM IN 5 YEARS.
X(X-10)= PRODUCT.